OpenStudy (goformit100):

You need to find the age of a women. Take double the age and add 4 more than the square root of twice the age. This sum added to its square will be equal to 2162.

OpenStudy (goformit100):

@ajprincess

OpenStudy (goformit100):

@TessaMarie08181997 @mikaela19900630

OpenStudy (ajprincess):

Let us assume the age of the woman be x. then double the age will be 2x. 4 more than the square root of twice the age is \(4+\sqrt{2x}\). their sum is \(2x+4+\sqrt{2x}\). the square of it will be \((2x+4+\sqrt{2x})^2\). their sum is \(2x+4+\sqrt{2x}+(2x+4+\sqrt{2x})^2=2162\). I guess this is the equation u need to solve.

OpenStudy (goformit100):

Oh Princess Thanks A lot . You are so Beautiful. ;) jk

OpenStudy (ajprincess):

lol welcome:)